You are evaluating two different cookie-baking ovens. The Pillsbury 707 costs $72,000, has a 5-year life, and has an annual OCF (after-tax) of –$11,700 per year. The Keebler CookieMunster costs $98,500, has a 7-year life, and has an annual OCF (after-tax) of –$9,700 per year.

If your discount rate is 13 percent, what is each machine’s EAC?
Answer #1

We use the concept of Equivalent Annual Cost (EAC), when the useful life of two alternatives are different.

Pillsbury 707 Keebler CookieMunster
a Purchase Cost ($)                   72,000                                         98,500
b Life of machines (years)                              5                                                   7
c Running cost per year ($)                   11,700                                           9,700
d PVAF (based on life and 13% discount factor)#                      3.517                                           4.423
e Present Value of Running cost of machine (c*d)                   41,152                                         42,899
f Cash outflow of machines (a+e)                 113,152                                      141,399
g Equivalent Annual Cost (f/d)                   32,171                                         31,972

Decision: Here EAC is low for Keebler CookieMunster, it should be selected.

#You can use the equation [(1-(1+r)-n)/r] to find PVAF using calculator.

PVAF13%,5 years = (1-1.13^-5)/0.13 = 3.5172312615427

PVAF13%,7 years = (1-1.13^-7)/0.13 = 4.42261043272199